# Large Language Models for Data Annotation: A Survey

### v2

@@ 4.2 Filtering & Selection
Rule-Based Methods. Rule-based methods follow various heuristic assumptions concerning sample length Li et al. (2023f); Kim et al. (2023a), keyword occurrence Kim et al. (2023b); Zheng et al. (2023) and specific patterns Zhang and Yang (2023a); Guo et al. (2024a); Ding et al. (2024) to filter low-quality or undesiered synthetic data points. Zheng et al. (2023); Kim et al. (2023a) establish thresholds for the number of rounds in generated conversations to guarantee each synthetic dialogue is informative enough. Ho et al. (2023); Kang et al. (2024) employ ground truth parsing to filter out incorrect CoT rationales within each candidate reasoning sample. To encourage diversity among the generated data points, Wang et al. (2022b); Lee et al. (2023a); Ding et al. (2024) utilize semantic similarity metrics to identify and remove redundant samples.
External-Source-Based Methods. There are also many works that depend on the external source’s feedback to clean and refine synthetic datasets Kim et al. (2023a). With a pre-trained reward model, Gulcehre et al. (2023); Dong et al. (2023) augment the original dataset only with samples that obtain high reward values. When distilling smaller models, Lin et al. (2023b); Wang et al. (2024d) meticulously select appropriate data through the feedback from the student models. Other approaches Chen et al. (2023g); Zheng et al. (2023) utilize pre-trained classification models to discern between target and unwanted data points.
LLMs-Driven Methods. The versatility of LLMs has invoked interest in leveraging LLMs themselves to do data selection. Some approaches use signals or features produced by LLMs, such as perplexity score Wang et al. (2024c), confidence levels Wang et al. (2022b); Huang et al. (2023), and logits Pace et al. (2024), as criteria for constructing data selectors. Others directly prompt the LLMs for this task. For instance, Lu et al. (2023) query the target LLM to assess the quality of generated samples. Kim et al. (2023a) leverage ChatGPT to determine if the social commonsense knowledge is appropriately conveyed in the synthetic dialogues. Additionally, there are also works that adopt the LLMs to rank multiple candidate annotations and utilize the top ones in the subsequent stages Jeronymo et al. (2023); Li et al. (2024c). In pairwise feedback synthesis, Tong et al. (2024b) task the base LLM with judging whether one response genuinely surpasses another.
@@ 5.1 Supervised Fine-tuning
Self-distillation. Huang et al. (2023) first propose the concept of self-improve that utilizes LLMs as both data annotators and learnable models and iteratively fine-tune LLMs in their self-annotated data. Wang et al. (2023e) also tune a GPT3 in the instruction tuning dataset to improve its zero-shot generalization capability. To foster LLMs’ evolution, Lu et al. (2023) iteratively fine-tune the LLMs in self-refined synthetic responses. To mitigate the distribution gap between task datasets and the LLMs, Yang et al. (2024b) use self-distillation which guides fine-tuning with a distilled dataset generated by the model itself. Both Chen et al. (2024b) and Cheng et al. (2024) introduce a self-play mechanism, where the LLM refines its capability by playing against instances of itself.
Distill Smaller Models. For efficiency issues, many studies aim to use the data generated by a large and powerful LLM to train a flexible and affordable smaller model. For a better instruction-following ability, many medium and small-sized LLMs are trained on the synthetic dataset produced by larger LLMs Taori et al. (2023); Chiang et al. (2023b); Xu et al. (2023a). In classification tasks, Meng et al. (2022, 2023); Wang et al. (2023d) augment the original datasets and train smaller bidirectional attention models on them. To foster models’ reasoning ability, many studies tune smaller models with synthetic rationales collected from LLMs Wang et al. (2022a); Shridhar et al. (2023); Liu et al. (2023a); Kang et al. (2024). Other task-specific capabilities distillation from LLMs include dialogue generation Xu et al. (2023b), information extraction Josifoski et al. (2023); Jeronymo et al. (2023) and code generation Chaudhary (2023); Roziere et al. (2023). Moreover, LLMs have been proven to follow a scaling law in terms of their knowledge capacity. Therefore, there is also a growing interest in distilling vertical and domain-specific knowledge from LLMs, including medicine Zhang et al. (2023); Xiong et al. (2023), finance Zhang and Yang (2023b) and science Luo et al. (2023); Zhao et al. (2024), to smaller models.
@@ 5.2 Alignment Tuning
Reward Modeling. LLMs-generated annotations can be used to train or refine the reward model for better alignment. Xu et al. (2023c) propose a data curriculum method that leverages the pairwise feedback from LLMs to calculate the sample difficulty level and smooth LLMs’ learning from simple ones to hard ones. Kim et al. (2023b) design reward model guided self-play to iteratively improve the reward model with synthesized data generated by the policy model. Pace et al. (2024) propose to maximize the probability of correctly labeling a pair of on-policy responses to a given query according to the base preference model. In robotics, Zeng et al. (2024) learns a reward function from scratch using the LLMs’ feedback. With synthetic data pair, Sun et al. (2024a) train an instructable reward model to generate reward scores based on arbitrary human-defined principles.
Policy Training. While many direct alignment methods Rafailov et al. (2024); Zhao et al. (2023b) have emerged recently, some works directly explore the use of annotated feedback for policy training. One common strategy is to directly apply DPO with the synthetic pairwise feedback produced by LLMs Yuan et al. (2024); Zhang et al. (2024a); Lee et al. (2024b); Tong et al. (2024b); Lee et al. (2024a); Guo et al. (2024b). Besides, Gulcehre et al. (2023); Dong et al. (2023) leverage a pre-trained reward model to filter low-quality synthetic data and iteratively tune LLMs with growing datasets. Wang et al. (2024a) propose a bootstrapping self-alignment method to repeatly utilize the synthetic data. Liu et al. (2024) introduce the Mixture of insighTful Experts (MoTE) architecture, which applies the mixture of experts to enhance each component of the synthetic response, markedly increasing alignment efficiency. With the reasoning pairwise feedback generated by LLM itself, Pang et al. (2024a) use a modified DPO loss with an additional negative log-likelihood term to tune the LLM.
@@ 5.3 Inference
In-context Learning. In-context Learning (ICL) consists of three components: a task description (or prompt), several in-context samples (or demonstration), and the test case that needs to be inferred. Current studies have applied the annotations and data generated by LLMs in all these components for refining or augmenting. Zhou et al. (2022b) first showed that with a well-designed pipeline, LLMs can be human-level prompt engineers to generate accurate task descriptions. Following them,  Yang et al. (2023b); Li et al. (2024e) conduct augmentation and expansion to the original task prompt, making it more detailed for LLMs to follow. Demonstration augmentation Kim et al. (2022); Li et al. (2023c); Chen et al. (2023d); He et al. (2024) is another useful skill to enrich and diversify the provided demonstrations, especially when the labeled data is limited. For the test sample, one augmentation method is to leverage LLMs to rephrase it once Deng et al. (2023) or multiple times Li et al. (2023a); Yang et al. (2024a). Other works study how to polish the original test sample Xi et al. (2023) or decompose it into several sub-questions Wang et al. (2024b).
@@ 6.1 Ethics Consideration
One critical concern of LLM-generated annotations is the ethics consideration, especially in high-stakes decision-making tasks like finance Yang et al. (2023a), jurisprudence Cui et al. (2023), and healthcare Eloundou et al. (2023). Despite the efficiency of LLM annotation, the lack of human insight may lead to biased and unfair results Wu et al. (2023); Abid et al. (2021); Cheng et al. (2021); Li et al. (2023g). Moreover, LLMs make human annotator roles redundant, potentially increasing social disparities Dillion et al. (2023). Future studies should harmonize technological advancements with societal consequences, including considering social implications, ensuring ethical use, promoting fairness, and maintaining transparency.
@@ 6.2 Challenges and Future Work
Model Collapse. Model collapse refers to the gradual performance decrease of an LLM trained on the outputs of other LLMs Sun et al. (2023); Gunasekar et al. (2023); Hsieh et al. (2023); Honovich et al. (2022); Chiang et al. (2023a); Geng et al. (2023). It is unavoidable since LLM-generated data is occupying the information ecosystem. The imitation model often replicates stylistic elements without achieving the factual precision of superior models Gudibande et al. (2023); Shumailov et al. (2023). This divergence is caused by statistical approximation error from limited sample sizes and functional approximation error from constrained model capacity. Both errors tend to amplify through successive training cycles Alemohammad et al. (2023).
@@
Hallucinations. Hallucinations in LLMs significantly undermine the integrity and reliability of their generated annotations Alkaissi and McFarlane (2023); Azamfirei et al. (2023); Chaudhary et al. (2024). Hullicinated outputs detached from factual information can cause the proliferation of misinformation Jiang et al. (2024); Chen and Shu (2023). Addressing hallucinations requires refining the training process and implementing validation mechanisms for annotations through automated and manual verification Liao and Vaughan (2023); Pan et al. (2023); Bian et al. (2023). Moreover, the inherent opacity of LLMs complicates efforts to investigate the causes of hallucinations.
Potential Solution. Yang et al. (2023d) addresses hallucinations in LLMs with the Reverse Validation method, detecting hallucinations at the passage level by constructing a query from the response and checking for a match within the LLM’s internal knowledge. Bertaglia et al. (2023) uses Chain-of-Thought (CoT) prompting and explanation generation, where CoT prompting produces explanations for predictions, ensuring logical and verifiable outputs. Li et al. (2023b) proposes the CoAnnotating framework, which uses uncertainty-guided work allocation between humans and LLMs, applying self-evaluation and entropy metrics to assess reliability and distribute tasks effectively.
@@ 7 Conclusion
The exploration of LLMs for data annotation has revealed an exciting frontier in NLP, presenting novel solutions to longstanding challenges like data scarcity, and enhancing annotation quality and process efficiency. This survey meticulously reviews methodologies, applications, and hurdles associated with LLM employment, including detailed taxonomy from annotation generation to utilization. It evaluates the effects of LLM-generated annotations on training machine learning models while addressing both technical and ethical concerns like bias and societal ramifications. Highlighting our novel taxonomy of LLM methodologies, strategies for utilizing LLM-generated annotations, and a critical discussion on the challenges, this work aims to steer future progress in this crucial area. Additionally, we introduce a comprehensive categorization of techniques and compile extensive benchmark datasets to support ongoing research endeavors, concluding with an examination of persistent challenges and open questions, paving the way for future investigative pursuits in the domain.
@@ 16 Nvidia V100 GPUs
  • Note: [1]Yoo et al. (2021); [2]Wang et al. (2023e); [3]Meng et al. (2023); [4]Wang et al. (2024c); [5]Li et al. (2023a); [6]Köksal et al. (2024); [7]Yu et al. (2024); [8]Zhang and Yang (2023a); [9]Huang et al. (2023); [10]Yang et al. (2024b); [11]Liu et al. (2024); [12]Guo et al. (2024a); [13]Sun et al. (2024b); [14]Wang et al. (2024a); [15]Luo et al. (2024); [16]Kojima et al. (2022); [17]Yao et al. (2024); [18]Hao et al. (2023); [19]Besta et al. (2024); [20]Yao et al. (2023); [21]Wang et al. (2024e); [22]Chen et al. (2023e); [23]Qi et al. (2023); [24]Tan et al. (2023); [25]Wang et al. (2022a); [26]Wang et al. (2023b); [27]Liu et al. (2023a); [28]Shridhar et al. (2023); [29]Kang et al. (2024); [30]Gao et al. (2021); [31]Wang et al. (2022b); [32]Chen et al. (2023f); [33]Liu et al. (2023c); [34]Tong et al. (2023); [35]Balepur et al. (2023); [36]Ma and Du (2023); [37]Yin et al. (2023); [38]Liang et al. (2023); [39]Xu et al. (2023d); [401]Liu et al. (2023e); [41]Bai et al. (2022); [42]Lee et al. (2023b); [43]Yuan et al. (2024); [44]Sun et al. (2024a); [45]Zhang et al. (2024a); [46]Pace et al. (2024); [47]Zeng et al. (2024); [48]Kim et al. (2023b); [49]Tong et al. (2024b); [50]Yang et al. (2023c); [51]Xu et al. (2023c); [52]Lee et al. (2024a); [53]Feng et al. (2024).
@@
  • Note: [1]Madaan et al. (2024); [2]Shinn et al. (2024); [3]Chen et al. (2023c); [4]Raunak et al. (2023); [5]Yang et al. (2023d); [6]Manakul et al. (2023); [7]Du et al. (2023a); [8]Xu et al. (2023d); [9]Cohen et al. (2023); [10]Fu et al. (2023); [11]Li et al. (2023d); [12]Chu et al. (2024b); [13]Ning et al. (2024); [14]Wang and Li (2023); [15]An et al. (2023); [16]Chen et al. (2023a); [17]Tong et al. (2024a); [18]Kim et al. (2023a); [19]Xu et al. (2023b); [20]Chen et al. (2023b); [21]Li et al. (2024d); [22]Zheng et al. (2023); [23]Chen et al. (2022); [24]Zhou et al. (2022a); [25]Xiang et al. (2022); [26]Li et al. (2024b); [27]Ronzano and Nanavati (2024); [28]Li et al. (2024c); [29]Ding et al. (2024); [30]Xiong et al. (2024); [31]Tuozzo (2022); [32]Huang et al. (2022); [33]Brohan et al. (2023); [34]Rana et al. (2023); [35]Singh et al. (2023); [36]Lin et al. (2023a); [37]Wang et al. (2023a); [38]Ha et al. (2023); [39]Kwon et al. (2022); [40]Du et al. (2023b); [41]Li et al. (2023f); [42]Yin et al. (2024); [43]Chen et al. (2024a); [44]Acharya et al. (2023); [45]Shen et al. (2024); [46]Wei et al. (2024); [47]Zhang et al. (2024b); [48]Josifoski et al. (2023); [49]Jeronymo et al. (2023); [50]Li et al. (2024a); [51]Ma et al. (2024); [52]Bonn et al. (2024); [53Chu et al. (2024a); [54]Bhattacharjee et al. (2024); [55]Martorana et al. (2024).
@@
  • Note: [1]Bai et al. (2022); [2]Kim et al. (2023a); [3]Kim et al. (2023b); [4]Zheng et al. (2023); [5]Zhang and Yang (2023a); [6]Guo et al. (2024a); [7]Ding et al. (2024); [8]Ho et al. (2023); [9]Kang et al. (2024); [10]Wang et al. (2022b); [11]Lee et al. (2023a); [12]Ding et al. (2024); [13]Gulcehre et al. (2023); [14]Dong et al. (2023); [15]Lin et al. (2023b); [16]Wang et al. (2024d); [17]Chen et al. (2023g); [18]Huang et al. (2023); [19]Pace et al. (2024); [20]Lu et al. (2023); [21]Jeronymo et al. (2023); [22]Li et al. (2024c); [23]Tong et al. (2024b).
@@
  • Note: [1]Huang et al. (2023); [2]Wang et al. (2023e); [3]Lu et al. (2023); [4]Yang et al. (2024b); [5]Chen et al. (2024b); [6]Cheng et al. (2024); [7]Taori et al. (2023); [8]Chiang et al. (2023a); [9]Xu et al. (2023a); [10]Meng et al. (2022); [11]Meng et al. (2023); [12]Wang et al. (2023d); [13]Wang et al. (2022a); [14]Shridhar et al. (2023); [15]Liu et al. (2023a); [16]Kang et al. (2024); [17]Xu et al. (2023b); [18]Josifoski et al. (2023); [19]Jeronymo et al. (2023); [20]Chaudhary (2023); [21]Roziere et al. (2023); [22]Zhang et al. (2023); [23]Xiong et al. (2023); [24]Zhang and Yang (2023b); [25]Luo et al. (2023); [26]Zhao et al. (2024); [27]Xu et al. (2023c); [28]Kim et al. (2023b); [29]Pace et al. (2024); [30]Zeng et al. (2024); [31]Sun et al. (2024a); [32]Yuan et al. (2024); [33]Zhang et al. (2024a); [34]Lee et al. (2024b); [35]Tong et al. (2024b); [36]Lee et al. (2024a); [37]Guo et al. (2024b); [38]Gulcehre et al. (2023); [39]Dong et al. (2023); [40]Wang et al. (2024a); [41]Liu et al. (2024); [42]Chen et al. (2023c); [43]Zhou et al. (2022b); [44]Yang et al. (2023b); [45]Li et al. (2024e); [46]Kim et al. (2022); [47]Li et al. (2023c); [48]Chen et al. (2023d); [49]He et al. (2024); [50]Deng et al. (2023); [51]Li et al. (2023a); [52]Yang et al. (2024a); [53]Xi et al. (2023); [54]Wang et al. (2024b); [55]Kojima et al. (2022); [56]Wang et al. (2022b); [57]Chen et al. (2023f); [58]Tong et al. (2023); [59]Balepur et al. (2023); [60]Ma and Du (2023); [61]Madaan et al. (2024); [62]Tong et al. (2024a); [63]Chen et al. (2023e); [64]Besta et al. (2024); [65]Hao et al. (2023).

### v3

@@ 4.1 Evaluating LLM-Generated Annotations
LLM-as-a-Judge: LLM-as-a-judge Wu et al. (2024c); Zheng et al. (2023b); Li et al. (2024c) is a commonly used method in automatic generation evaluation. To scale the assessment of the synthetic data or annotation, there are also some works that adopt LLM-as-a-judge to conduct the evaluation. Li et al. (2024g) employ multiple LLMs to debate with each other to evaluate the synthetic data’s quality fairly, iteratively improving response quality, while creating a judge LLM to select preferred responses for enhanced instruction tuning. To enhance the quality of the synthetic instruction tuning data, Liang et al. (2024b) introduce an iterative self-enhancement paradigm (I-SHEEP). During training, they adopt LLM-as-a-judge to score the synthetic responses and set a threshold to collect high-quality query-response pairs for the subsequent training iteration.
@@ 4.2 Filtering & Selection
Rule-Based Methods. Rule-based methods follow various heuristic assumptions concerning sample length Li et al. (2023f); Kim et al. (2023a), keyword occurrence Kim et al. (2023b); Zheng et al. (2023a) and specific patterns Zhang and Yang (2023a); Guo et al. (2024a); Ding et al. (2024) to filter low-quality or undesiered synthetic data points. Zheng et al. (2023a); Kim et al. (2023a) establish thresholds for the number of rounds in generated conversations to guarantee each synthetic dialogue is informative enough. Ho et al. (2023); Kang et al. (2024) employ ground truth parsing to filter out incorrect CoT rationales within each candidate reasoning sample. To encourage diversity among the generated data points, Wang et al. (2022b); Lee et al. (2023a); Ding et al. (2024) utilize semantic similarity metrics to identify and remove redundant samples.
External-Source-Based Methods. There are also many works that depend on the external source’s feedback to clean and refine synthetic datasets Kim et al. (2023a). With a pre-trained reward model, Gulcehre et al. (2023); Dong et al. (2023) augment the original dataset only with samples that obtain high reward values. When distilling smaller models, Lin et al. (2023b); Wang et al. (2024f) meticulously select appropriate data through the feedback from the student models. Other approaches Chen et al. (2023g); Zheng et al. (2023a) utilize pre-trained classification models to discern between target and unwanted data points.
LLMs-Driven Methods. The versatility of LLMs has invoked interest in leveraging LLMs themselves to do data selection. Some approaches use signals or features produced by LLMs, such as perplexity score Wang et al. (2023f), confidence levels Wang et al. (2022b); Huang et al. (2023), and logits Pace et al. (2024), as criteria for constructing data selectors. Others directly prompt the LLMs for this task. For instance, Lu et al. (2023) query the target LLM to assess the quality of generated samples. Kim et al. (2023a) leverage ChatGPT to determine if the social commonsense knowledge is appropriately conveyed in the synthetic dialogues. Additionally, there are also works that adopt the LLMs to rank multiple candidate annotations and utilize the top ones in the subsequent stages Jeronymo et al. (2023); Li et al. (2024e). In pairwise feedback synthesis, Tong et al. (2024b) task the base LLM with judging whether one response genuinely surpasses another. Besides, Jiang et al. (2024b) demonstrate that filtering out correct but with high distribution shift extent (DSE) samples could also benefit the results of self-improvement.
@@ 5.1 Supervised Fine-Tuning
Self-Evolution. Huang et al. (2023) first propose the concept of self-improve that utilizes LLMs as both data annotators and learnable models and iteratively fine-tune LLMs in their self-annotated data. Wang et al. (2023e) also tune a GPT3 in the instruction tuning dataset to improve its zero-shot generalization capability. To foster LLMs’ evolution, Lu et al. (2023) iteratively fine-tune the LLMs in self-refined synthetic responses. To mitigate the distribution gap between task datasets and the LLMs, Yang et al. (2024b) use self-distillation which guides fine-tuning with a distilled dataset generated by the model itself. Both Chen et al. (2024d) and Cheng et al. (2024b) introduce a self-play mechanism, where the LLM refines its capability by playing against instances of itself. Moreover, Wang et al. (2024b) demonstrate that the reasoning abilities of small-scale LMs can be enhanced through self-training, a process where models learn from their own outputs.
Distill Smaller Models. For efficiency issues, many studies aim to use the data generated by a large and powerful LLM to train a flexible and affordable smaller model. For a better instruction-following ability, many medium and small-sized LLMs are trained on the synthetic dataset produced by larger LLMs Taori et al. (2023); Chiang et al. (2023b); Xu et al. (2023a). In classification tasks, Meng et al. (2022, 2023); Wang et al. (2023d) augment the original datasets and train smaller bidirectional attention models on them. To foster models’ reasoning ability, many studies tune smaller models with synthetic rationales collected from LLMs Wang et al. (2022a); Shridhar et al. (2023); Liu et al. (2023a); Kang et al. (2024). Other task-specific capabilities distillation from LLMs include dialogue generation Xu et al. (2023b), information extraction Josifoski et al. (2023); Jeronymo et al. (2023) and code generation Chaudhary (2023); Roziere et al. (2023). Moreover, LLMs have been proven to follow a scaling law in terms of their knowledge capacity. Therefore, there is also a growing interest in distilling vertical and domain-specific knowledge from LLMs, including medicine Zhang et al. (2023); Xiong et al. (2023a), finance Zhang and Yang (2023b) and science Luo et al. (2023); Zhao et al. (2024), to smaller models.
@@ 5.2 Alignment Tuning
Reward Modeling. LLMs-generated annotations can be used to train or refine the reward model for better alignment. Xu et al. (2023c) propose a data curriculum method that leverages the pairwise feedback from LLMs to calculate the sample difficulty level and smooth LLMs’ learning from simple ones to hard ones. Kim et al. (2023b) design reward model guided self-play to iteratively improve the reward model with synthesized data generated by the policy model. Pace et al. (2024) propose to maximize the probability of correctly labeling a pair of on-policy responses to a given query according to the base preference model. In robotics, Zeng et al. (2024) learns a reward function from scratch using the LLMs’ feedback. With synthetic data pair, Sun et al. (2023b) train an instructable reward model to generate reward scores based on arbitrary human-defined principles.
Policy Training. While many direct alignment methods Rafailov et al. (2024); Zhao et al. (2023b) have emerged recently, some works directly explore the use of annotated feedback for policy training. One common strategy is to directly apply DPO with the synthetic pairwise feedback produced by LLMs Yuan et al. (2024); Zhang et al. (2024b); Lee et al. (2024b); Tong et al. (2024b); Lee et al. (2024a); Guo et al. (2024b). Besides, Gulcehre et al. (2023); Dong et al. (2023) leverage a pre-trained reward model to filter low-quality synthetic data and iteratively tune LLMs with growing datasets. Wang et al. (2024a) propose a bootstrapping self-alignment method to repeatly utilize the synthetic data. Liu et al. (2024c) introduce the Mixture of insighTful Experts (MoTE) architecture, which applies the mixture of experts to enhance each component of the synthetic response, markedly increasing alignment efficiency. With the reasoning pairwise feedback generated by LLM itself, Pang et al. (2024a) use a modified DPO loss with an additional negative log-likelihood term to tune the LLM.
@@ 5.3 Inference
In-Context Learning. In-context Learning (ICL) consists of three components: a task description (or prompt), several in-context samples (or demonstration), and the test case that needs to be inferred. Current studies have applied the annotations and data generated by LLMs in all these components for refining or augmenting. Zhou et al. (2022b) first showed that with a well-designed pipeline, LLMs can be human-level prompt engineers to generate accurate task descriptions. Following them,  Yang et al. (2023b); Li et al. conduct augmentation and expansion to the original task prompt, making it more detailed for LLMs to follow. Demonstration augmentation Kim et al. (2022); Li et al. (2023c); Chen et al. (2023d); He et al. (2024) is another useful skill to enrich and diversify the provided demonstrations, especially when the labeled data is limited. For the test sample, one augmentation method is to leverage LLMs to rephrase it once Deng et al. (2023) or multiple times Li et al. (2023a); Yang et al. (2024a). Other works study how to polish the original test sample Xi et al. (2023) or decompose it into several sub-questions Wang et al. (2024b).
@@ 6.1 Ethics Consideration
One critical concern of LLM-generated annotations is the ethics consideration, especially in high-stakes decision-making tasks like finance Yang et al. (2023a), jurisprudence Cui et al. (2023), and healthcare Eloundou et al. (2023). Despite the efficiency of LLM annotation, the lack of human insight may lead to biased and unfair results Wu et al. (2023); Abid et al. (2021); Cheng et al. (2021); Li et al. (2023g); Beigi et al. (2024); Das et al. (2024); Shimabucoro et al. (2024). Moreover, LLMs make human annotator roles redundant, potentially increasing social disparities Dillion et al. (2023). Future studies should harmonize technological advancements with societal consequences, including considering social implications, ensuring ethical use, promoting fairness, and maintaining transparency.
@@ 6.2 Challenges and Future Work
Model Collapse. Model collapse refers to the gradual performance decrease of an LLM trained on the outputs of other LLMs Sun et al. (2023a); Gunasekar et al. (2023); Hsieh et al. (2023); Honovich et al. (2022); Chiang et al. (2023a); Geng et al. (2023); Huang et al. (2024a). It is unavoidable since LLM-generated data is occupying the information ecosystem. The imitation model often replicates stylistic elements without achieving the factual precision of superior models Gudibande et al. (2023); Shumailov et al. (2023). This divergence is caused by statistical approximation error from limited sample sizes and functional approximation error from constrained model capacity. Both errors tend to amplify through successive training cycles Alemohammad et al. (2023).
@@
Hallucinations. Hallucinations in LLMs significantly undermine the integrity and reliability of their generated annotations Alkaissi and McFarlane (2023); Azamfirei et al. (2023); Chaudhary et al. (2024). Hullicinated outputs detached from factual information can cause the proliferation of misinformation Jiang et al. (2024a); Chen and Shu (2023, ); Huang et al. (2024b). Addressing hallucinations requires refining the training process and implementing validation mechanisms for annotations through automated and manual verification Liao and Vaughan (2023); Pan et al. (2023); Bian et al. (2023). Moreover, the inherent opacity of LLMs complicates efforts to investigate the causes of hallucinations.
Potential Solution. Yang et al. (2023d) addresses hallucinations in LLMs with the Reverse Validation method, detecting hallucinations at the passage level by constructing a query from the response and checking for a match within the LLM’s internal knowledge. Bertaglia et al. (2023) uses Chain-of-Thought (CoT) prompting and explanation generation, where CoT prompting produces explanations for predictions, ensuring logical and verifiable outputs. Li et al. (2023b) proposes the CoAnnotating framework, which uses uncertainty-guided work allocation between humans and LLMs, applying self-evaluation and entropy metrics to assess reliability and distribute tasks effectively. Zendel et al. (2024) propose a human-LLM connotation process for better annotation quality.
@@ 7 Conclusion
The exploration of LLMs for data annotation and synthesis has revealed an exciting frontier in NLP, presenting novel solutions to longstanding challenges like data scarcity, and enhancing annotation quality and process efficiency. This survey meticulously reviews methodologies, applications, and hurdles associated with LLM employment, including detailed taxonomy from annotation generation to utilization. It evaluates the effects of LLM-generated annotations on training machine learning models while addressing both technical and ethical concerns like bias and societal ramifications. Highlighting our novel taxonomy of LLM methodologies, strategies for utilizing LLM-generated annotations, and a critical discussion on the challenges, this work aims to steer future progress in this crucial area. Additionally, we introduce a comprehensive categorization of techniques and compile extensive benchmark datasets to support ongoing research endeavors, concluding with an examination of persistent challenges and open questions, paving the way for future investigative pursuits in the domain.
@@
Acknowledgements
The material in this presentation is supported by the National Science Foundation (NSF) under grants IIS-2229461, and the U.S. Department of Homeland Security under Grant Award Number, 17STQAC00001-08-00 and the U.S. Office of Naval Research (ONR) under grant N00014-21-1-4002. Lu Cheng is supported by the National Science Foundation (NSF) Grant #2312862, NIH #R01AG091762, and a Cisco gift grant. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Department of Homeland Security and the National Science Foundation.
@@ 16 Nvidia V100 GPUs
  • Note: [1]Yoo et al. (2021); [2]Wang et al. (2023e); [3]Meng et al. (2023); [4]Wang et al. (2023f); [5]Li et al. (2023a); [6]Köksal et al. ; [7]Yu et al. (2024); [8]Zhang and Yang (2023a); [9]Huang et al. (2023); [10]Yang et al. (2024b); [11]Liu et al. (2024c); [12]Guo et al. (2024a); [13]Sun et al. (2024b); [14]Wang et al. (2024a); [15]Luo et al. (2024a); [16]Kojima et al. (2022); [17]Yao et al. (2024); [18]Hao et al. (2023); [19]Besta et al. (2024); [20]Yao et al. (2023); [21]Wang et al. (2024g); [22]Chen et al. (2023e); [23]Qi et al. (2023); [24]Tan et al. (2023); [25]Wang et al. (2022a); [26]Wang et al. (2023b); [27]Liu et al. (2023a); [28]Shridhar et al. (2023); [29]Kang et al. (2024); [30]Gao et al. (2021); [31]Wang et al. (2022b); [32]Chen et al. (2023f); [33]Liu et al. (2023c); [34]Tong et al. (2023); [35]Balepur et al. (2023); [36]Ma and Du (2023); [37]Yin et al. (2023); [38]Liang et al. (2023); [39]Xu et al. (2023d); [401]Liu et al. (2023e); [41]Bai et al. (2022); [42]Lee et al. (2023b); [43]Yuan et al. (2024); [44]Sun et al. (2023b); [45]Zhang et al. (2024b); [46]Pace et al. (2024); [47]Zeng et al. (2024); [48]Kim et al. (2023b); [49]Tong et al. (2024b); [50]Yang et al. (2023c); [51]Xu et al. (2023c); [52]Lee et al. (2024a); [53]Feng et al. (2024).
@@
  • Note: [1]Madaan et al. (2024); [2]Shinn et al. (2024); [3]Chen et al. (2023c); [4]Raunak et al. (2023); [5]Yang et al. (2023d); [6]Manakul et al. (2023); [7]Du et al. (2023a); [8]Xu et al. (2023d); [9]Cohen et al. (2023); [10]Fu et al. (2023); [11]Li et al. (2023d); [12]Chu et al. (2024b); [13]Ning et al. (2024); [14]Wang and Li (2023); [15]An et al. (2023); [16]Chen et al. (2023a); [17]Tong et al. (2024a); [18]Kim et al. (2023a); [19]Xu et al. (2023b); [20]Chen et al. (2023b); [21]Li et al. (2024f); [22]Zheng et al. (2023a); [23]Chen et al. (2022); [24]Zhou et al. (2022a); [25]Xiang et al. (2022); [26]Li et al. (2024d); [27]Ronzano and Nanavati (2024); [28]Li et al. (2024e); [29]Ding et al. (2024); [30]Xiong et al. (2024a); [31]Tuozzo (2022); [32]Huang et al. (2022); [33]Brohan et al. (2023); [34]Rana et al. (2023); [35]Singh et al. (2023); [36]Lin et al. (2023a); [37]Wang et al. (2023a); [38]Ha et al. (2023); [39]Kwon et al. (2022); [40]Du et al. (2023b); [41]Li et al. (2023f); [42]Yin et al. (2024); [43]Chen et al. (2024c); [44]Acharya et al. (2023); [45]Shen et al. (2024); [46]Wei et al. (2024a); [47]Zhang et al. (2024c); [48]Josifoski et al. (2023); [49]Jeronymo et al. (2023); [50]Li et al. (2024b); [51]Ma et al. (2024); [52]Bonn et al. (2024); [53Chu et al. (2024a); [54]Bhattacharjee et al. (2024); [55]Martorana et al. (2024a).
@@
  • Note: [1]Bai et al. (2022); [2]Kim et al. (2023a); [3]Kim et al. (2023b); [4]Zheng et al. (2023a); [5]Zhang and Yang (2023a); [6]Guo et al. (2024a); [7]Ding et al. (2024); [8]Ho et al. (2023); [9]Kang et al. (2024); [10]Wang et al. (2022b); [11]Lee et al. (2023a); [12]Ding et al. (2024); [13]Gulcehre et al. (2023); [14]Dong et al. (2023); [15]Lin et al. (2023b); [16]Wang et al. (2024f); [17]Chen et al. (2023g); [18]Huang et al. (2023); [19]Pace et al. (2024); [20]Lu et al. (2023); [21]Jeronymo et al. (2023); [22]Li et al. (2024e); [23]Tong et al. (2024b).
@@
  • Note: [1]Huang et al. (2023); [2]Wang et al. (2023e); [3]Lu et al. (2023); [4]Yang et al. (2024b); [5]Chen et al. (2024d); [6]Cheng et al. (2024b); [7]Taori et al. (2023); [8]Chiang et al. (2023a); [9]Xu et al. (2023a); [10]Meng et al. (2022); [11]Meng et al. (2023); [12]Wang et al. (2023d); [13]Wang et al. (2022a); [14]Shridhar et al. (2023); [15]Liu et al. (2023a); [16]Kang et al. (2024); [17]Xu et al. (2023b); [18]Josifoski et al. (2023); [19]Jeronymo et al. (2023); [20]Chaudhary (2023); [21]Roziere et al. (2023); [22]Zhang et al. (2023); [23]Xiong et al. (2023a); [24]Zhang and Yang (2023b); [25]Luo et al. (2023); [26]Zhao et al. (2024); [27]Xu et al. (2023c); [28]Kim et al. (2023b); [29]Pace et al. (2024); [30]Zeng et al. (2024); [31]Sun et al. (2023b); [32]Yuan et al. (2024); [33]Zhang et al. (2024b); [34]Lee et al. (2024b); [35]Tong et al. (2024b); [36]Lee et al. (2024a); [37]Guo et al. (2024b); [38]Gulcehre et al. (2023); [39]Dong et al. (2023); [40]Wang et al. (2024a); [41]Liu et al. (2024c); [42]Chen et al. (2023c); [43]Zhou et al. (2022b); [44]Yang et al. (2023b); [45]Li et al. ; [46]Kim et al. (2022); [47]Li et al. (2023c); [48]Chen et al. (2023d); [49]He et al. (2024); [50]Deng et al. (2023); [51]Li et al. (2023a); [52]Yang et al. (2024a); [53]Xi et al. (2023); [54]Wang et al. (2024b); [55]Kojima et al. (2022); [56]Wang et al. (2022b); [57]Chen et al. (2023f); [58]Tong et al. (2023); [59]Balepur et al. (2023); [60]Ma and Du (2023); [61]Madaan et al. (2024); [62]Tong et al. (2024a); [63]Chen et al. (2023e); [64]Besta et al. (2024); [65]Hao et al. (2023).